# Safe Pareto Improvements for Expected Utility Maximizers in Program Games

### v4

@@ 3.4 Incentives to renegotiate
  1. 1.
  2. 2.
@@
Let $G(\mathcal{P})$ be any program game. Let $\boldsymbol{\beta}$ be any belief profile satisfying the assumption that players are certain that renegotiation won’t be punished (Assumption 5). Then, for every subjective equilibrium $(\textbf{p}^{*},\boldsymbol{\beta})$ of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{rn}})$ where $p_{i}^{*}\notin P_{i}^{\textnormal{rn}}$ for some $i$, there exists $\textbf{p}^{\prime}$ such that:
  1. 1.
    For all $i$, we have $p_{i}^{\prime}\in P_{i}^{\textnormal{rn}}$.
  2. 2.
    $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{rn}})$.
@@
4 The SPI Selection Problem and Conditional Set-Valued Renegotiation
@@
Figure 1: Illustration of the Pareto meet projection (PMP) of three different outcomes (black points) in the Scheduling Game, for each player. Gray points represent payoffs at each pure strategy profile. Each black point is mapped via a player’s PMP (black arrows) to a set containing a) the “nearest” point in the Pareto meet and b) all points better for the given player and no better for the other player than (a).
Here, we argue that, although in part the players’ initial bargaining problem recurs in SPI selection, players will always renegotiate so that each attains at least the worst payoff they can get in any efficient outcome. Following Rabin (1991) we call the profile of these payoffs the Pareto meet minimum (PMM). Player $i$’s Pareto meet projection (PMP) (Fig. 1) maps each outcome to the set of Pareto improvements such that, first, each player’s payoff is at least the PMM, and second, the payoff of each $j\neq i$ is not increased except up to the PMM. We will prove our bound by arguing that if players attempt to negotiate a Pareto improvement on an outcome, they always at least weakly prefer to be willing to negotiate to the PMP of that outcome.
@@
Let ${E}$ be the set of Pareto-efficient action profiles in $G$. Then the Pareto meet minimum (PMM) payoff profile is $\textbf{u}^{\textnormal{PMM}}=(\min_{\boldsymbol{a}\in{E}}u_{i}(\boldsymbol{a}))_{i=1}^{{n}}$. Player $i$’s Pareto meet projection (PMP) of an action profile $\boldsymbol{a}$ is the set $\mathbf{PMP}_{i}(\boldsymbol{a})$ of action profiles $\tilde{\boldsymbol{a}}$ such that $u_{i}(\tilde{\boldsymbol{a}})\geq\max\{u_{i}^{\textnormal{PMM}},u_{i}(\boldsymbol{a})\}$ and for all $j\neq i$, we have $u_{j}(\tilde{\boldsymbol{a}})=\max\{u_{j}^{\textnormal{PMM}},u_{j}(\boldsymbol{a})\}$.
We’ll start by giving an informal description of the algorithm we will use to prove the guarantee, called conditional set-valued renegotiation (CSR). Next, we’ll describe different components of the algorithm in more depth. Then, we’ll formally present the algorithm and the guarantee. Finally, we’ll show that players are not always incentivized to use SPIs that improve on the PMM.
4.1 Overview of CSR
If players want to increase their chances of Pareto-improving via renegotiation, without necessarily accepting renegotiation outcomes that heavily favor their counterparts, they can report to each other multiple renegotiation outcomes they each would find acceptable and take Pareto improvements on which they agree. CSR implements such an approach. Like Algorithm 1, CSR involves default programs, and checks whether the default programs of a profile of CSR algorithms result in an efficient outcome. If not, CSR moves to a renegotiation procedure that works as follows:
  1. 1.
    Renegotiation using conditional sets. At this stage, programs “announce” sets of points that Pareto-improve on the default and that they are willing to renegotiate to, conditional on the other players’ programs. If these sets overlap, the procedure continues to the second step; otherwise the players revert to their defaults.
    The intuition for using sets at this stage is that (we will argue), if a player’s program would renegotiate to some point worse for them than the PMM, they can safely instead use a program that is also willing to renegotiate to a payoff above their PMM payoff — without risking miscoordination if the other players do not also choose this new payoff precisely. Renegotiation sets that condition on the other players’ renegotiation set functions are crucial to the result that players are guaranteed their PMM payoff. This is because unconditionally adding an outcome to the renegotiation set might provide Pareto improvements against some possible counterpart profile, but make the outcome worse against some other possible counterpart profile (see Example 4.2).
  2. 2.
@@ 4.1 Overview of CSR
4.2 Components of CSR
Set-valued renegotiation.
To avoid the need to coordinate on an exact renegotiation function, players can use functions that map miscoordinated outcomes to sets of Pareto improvements they each find acceptable. Then, we suppose the players follow some rule (a selection function) for choosing an efficient outcome from their agreement set.
@@ 4.2 Components of CSR
Let $\mathbf{C}(\mathcal{A})$ be the set of closed subsets of $\mathcal{A}$ [^8] I.e., closed with respect to the topology on $\mathcal{A}$ induced by the Euclidean distance $d(\boldsymbol{a},\boldsymbol{a}^{\prime})=||\textbf{u}(\boldsymbol{a})-\textbf{u}(\boldsymbol{a}^{\prime})||$. Letting $\mathcal{R}^{i}$ be a set of functions from ×$\bigtimes_{j\neq i}\mathcal{R}^{j}\times\mathcal{A}$ to $\mathbf{C}(\mathcal{A})$, a function $\mathbf{RN}^{i}\in\mathcal{R}^{i}$ is a set-valued renegotiation function if, for all $\mathbf{RN}^{-i}\in\bigtimes_{j\neq i}\mathcal{R}^{j}$:
  1. 1.
    For all $\boldsymbol{a}\in\mathcal{A}$ and $\boldsymbol{a}^{\prime}\in\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})$, we have $\textbf{u}(\boldsymbol{a}^{\prime})\succeq\textbf{u}(\boldsymbol{a})$.
  2. 2.
    For some $\boldsymbol{a}$ and some $\boldsymbol{a}^{\prime}\in\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})$, we have $u_{i^{\prime}}(\boldsymbol{a}^{\prime})>u_{i^{\prime}}(\boldsymbol{a})$ for some $i^{\prime}$.
@@
Note that there is no bargaining problem involved in coordinating on a selection function, which we take to justify our assumption of a fixed selection function. To see this, consider two players who intended to use renegotiation programs with different selection functions. Each player could switch to using a program that used the other player’s selection function, and modify their set-valued renegotiation function such that it was guaranteed to induce the same outcome as if the other player switched to their selection function. (See Appendix C for a formal argument.) So the players are indifferent as to which selection function is used. (Coordinating on a selection function is a pure coordination problem, however; compare to the problem of coordinating on the programming language used in syntactic comparison-based program equilibrium (Tennenholtz, 2004).) In the results that follow, we will show that players can guarantee the PMM no matter which (transitive) selection function they use.
@@
Figure 2: Two possible cases of set-valued renegotiation in the Scheduling Game, for different player 2 renegotiation sets. Black points represent renegotiation outcomes (mapped from the miscoordination outcome $(0,0)$). If player 1 uses the renegotiation set shown here, they can achieve a Pareto improvement even if players don’t reach the Pareto frontier (left), while still allowing for their best possible outcome (right).
@@
(Set-valued renegotiation) The two plots in Fig. 2 show two possible player 2 renegotiation sets for the Scheduling Game (Table 1), given that the players miscoordinate at (Slot 1, Slot 2), and the renegotiation outcomes achieved given player 1’s renegotiation set (black points). [^10] We assume players can use correlated randomization to play any convex combination of the base game outcomes. Player 1 thinks it’s likely that the only efficient outcome player 2 is willing to renegotiate to is their own most preferred outcome, (Slot 2, Slot 2) (left plot). But player 1 believes that with positive probability player 2’s renegotiation set will also include player 1’s most preferred outcome, (Slot 1, Slot 1) (right plot). Player 1’s best response given these beliefs may be to choose a set-valued renegotiation function that maps (Slot 1, Slot 2) to a set including both (Slot 1, Slot 1) and other outcomes Pareto-worse than (Slot 1, Slot 1), i.e., the set depicted in Fig. 2. This way, they still achieve a Pareto improvement if player 2 has the smaller set, and get their best payoff if player 2 has the larger set.
Conditional renegotiation sets.
If players don’t use conditional renegotiation sets, they may not be able to guarantee desired Pareto improvements against some players without risking worse outcomes against others. Suppose that for some $\boldsymbol{a}$, player $i$ considers using some $\mathbf{RN}^{i}$ such that $\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})$ does not depend on $\mathbf{RN}^{-i}$. Against some possible profile of counterparts, adding some outcome $\boldsymbol{a}^{*}$ to $i$’s renegotiation set would allow for a Pareto improvement on the renegotiation outcome they would have otherwise reached, without risking exploitation. However, $i$ worries that some other profile of counterparts might use renegotiation sets such that, for some selection function, the renegotiation outcome reached when using $\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})\cup\{\boldsymbol{a}^{*}\}$ is worse than if they had used $\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})$. So player $i$ prefers to use $\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a})$ overall.
@@
Figure 3: Two possible renegotiation procedures in the Scheduling Game, for different player 2 renegotiation sets. The star represents some point that player 1 might add to their unconditional renegotiation set. In the case in the left plot, player 1 is no worse off by adding the star point to their set. But in the case in the right plot, if player 1 adds the star point, they might do worse if the selection function chooses that point instead of the black point that would have otherwise been achieved.
@@
(Failure of unconditional renegotiation sets) Suppose that in the Scheduling Game, Player 1 uses an unconditional set-valued renegotiation function $\mathbf{RN}^{1}$. Fig. 3 shows their set for the miscoordination outcome (Slot 1, Slot 2) (same in both plots). Suppose player 1 instead considers using ${\mathbf{RN}^{1}}^{\prime}$ such that for all $\mathbf{RN}^{2}$, the set ${\mathbf{RN}^{1}}^{\prime}(\mathbf{RN}^{2},\boldsymbol{a})$ includes a Pareto improvement $\boldsymbol{a}^{*}$ on the black point in the left plot (star). Then, against a player 2 with the renegotiation set in the right plot, if the players renegotiate to $\boldsymbol{a}^{*}$, player 1 is worse off than otherwise. But if player 1 had access to conditional renegotiation sets, they could instead use an ${\mathbf{RN}^{1}}^{\prime}$ that maps (Slot 1, Slot 2) to $\boldsymbol{a}^{*}$ only against the player 2 set in the left plot.
Renegotiation sets that guarantee the PMM.

### v5

@@ 3.3. Incentives to Renegotiate
We say that players with beliefs $\boldsymbol{\beta}$ are certain that renegotiation won’t be punished if the following holds. Take any renegotiation function $\mathbf{rn}$ $\in\mathcal{R}$; any renegotiation program $p_{i}\in P_{i}^{\textnormal{rn}}(\mathbf{rn})$; and any $p_{j}$ in the support of $\beta_{i}$ such that the programs in p don’t renegotiate with each other. (I.e., there is no $\mathbf{rn}^{j}$ such that $p_{j}\in P_{j}^{\textnormal{rn}}(\mathbf{rn}^{j})$ where $\mathbf{rn}^{j}(\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))=\mathbf{rn}(\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$.) Then:
  1. (1)
    $p_{j}(p_{i})=p_{j}(p_{i}^{\textnormal{def}})$.
  2. (2)
    If $p_{i}^{\textnormal{def}}$ is used in subjective equilibrium with respect to $\beta_{i}$, and $p_{j}\in P_{j}^{\textnormal{rn}}$, we have $p_{i}^{\textnormal{def}}(p_{j})=p_{i}^{\textnormal{def}}(p_{j}^{\textnormal{def}})$.
@@
Let $G(\mathcal{P})$ be any program game. Let $\boldsymbol{\beta}$ be any belief profile satisfying the assumption that players are certain that renegotiation won’t be punished (Assumption 1). And, for some arbitrary renegotiation function $\mathbf{rn}$, for each $i$ and $p_{i}\in P_{i}$, let $f_{i}^{*}(p_{i})$ be the program of the form in Algorithm 1 with ${f_{i}^{*}(p_{i})}^{\textnormal{def}}=p_{i}$. Then, for every subjective equilibrium $(\textbf{p}^{*},\boldsymbol{\beta})$ of $G(\mathcal{P}\cup\mathcal{P}^{\text{rn}})$ where $p_{i^{\prime}}^{*}\notin P_{i^{\prime}}^{\textnormal{rn}}$ for some $i^{\prime}$, there exists $\textbf{p}^{\prime}\in\mathcal{P}^{\text{rn}}$ such that:
  1. (1)
    For all $i$,
    $p_{i}^{\prime}=\begin{cases}p_{i}^{*},&\text{if $p_{i}^{*}\in P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ for some $\mathbf{rn}^{i}$;}\\ f_{i}^{*}(p_{i}^{*}),&\text{else.}\end{cases}$
  2. (2)
    $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P}\cup\mathcal{P}^{\text{rn}})$.
@@
4. The SPI Selection Problem and Conditional Set-Valued Renegotiation
@@
Figure 1. Illustration of the Pareto meet projection (PMP) of three different outcomes (black points) in the Scheduling Game, for each player. Gray points represent payoffs at each pure strategy profile. Each black point is mapped via a player’s PMP (black arrows) to a set containing a) the “nearest” point in the Pareto meet and b) all points better for the given player and no better for the other player than (a).
[…]
This figure shows three different payoff plots in the Scheduling Game, each illustrating how the Pareto meet projection (PMP) maps different outcomes to sets of points. The graph shows payoff space with Player 1’s payoff on the x-axis and Player 2’s payoff on the y-axis. Gray points mark the payoffs from pure strategy profiles: (3,1), (1,3), and (1,1). The three black points represent different outcomes being mapped via PMP: (0,0), (1,1), and (2,2). Black arrows show how each point is mapped to its corresponding PMP set. For example, the miscoordination outcome (0,0) is mapped to a set that includes both the PMM point (1,1) and points that are better for the projecting player but no better for the other player.
Here, we argue that, although in part the players’ initial bargaining problem recurs in SPI selection, players will always renegotiate so that each attains at least the worst payoff they can get in any efficient outcome. Following Rabin (1991) we call the profile of these payoffs the Pareto meet minimum (PMM). Player $i$’s Pareto meet projection (PMP) (Fig. 1) maps each outcome to the set of Pareto improvements such that, first, each player’s payoff is at least the PMM, and second, the payoff of $j\neq i$ is not increased except up to the PMM. We will prove our bound by arguing that if players attempt to negotiate a Pareto improvement on an outcome, they always at least weakly prefer to be willing to negotiate to the PMP of that outcome.
@@
Let ${E}$ be the set of Pareto-efficient action profiles in $G$. Then the Pareto meet minimum (PMM) payoff profile is $\textbf{u}^{\textnormal{PMM}}=(\min_{\boldsymbol{a}\in{E}}u_{1}(\boldsymbol{a}),\min_{\boldsymbol{a}\in{E}}u_{2}(\boldsymbol{a}))$. Player $i$’s Pareto meet projection (PMP) of an action profile $\boldsymbol{a}$ is the set $\mathbf{PMP}_{i}(\boldsymbol{a})$ of action profiles $\tilde{\boldsymbol{a}}$ such that $u_{i}(\tilde{\boldsymbol{a}})\geq\max\{u_{i}^{\textnormal{PMM}},u_{i}(\boldsymbol{a})\}$ and $u_{j}(\tilde{\boldsymbol{a}})=\max\{u_{j}^{\textnormal{PMM}},u_{j}(\boldsymbol{a})\}$.
We’ll start by giving an informal description of the algorithm we will use to prove the guarantee, called conditional set-valued renegotiation (CSR). Next, we’ll describe different components of the algorithm in more depth. Finally, we’ll formally present the algorithm and the guarantee.
4.1. Overview of CSR
If players want to increase their chances of Pareto-improving via renegotiation, without necessarily accepting renegotiation outcomes that heavily favor their counterpart, they can report to each other multiple renegotiation outcomes they each would find acceptable and take Pareto improvements on which they agree. CSR implements such an approach. Like Algorithm 1, CSR involves default programs, and checks whether the default programs of a profile of CSR algorithms result in an efficient outcome. If not, CSR moves to a renegotiation procedure that works as follows:
  1. (1)
    Renegotiation using conditional sets. At this stage, programs “announce” sets of points that Pareto-improve on the default and that they are willing to renegotiate to, conditional on the other player’s program (see shaded regions in Fig. 2). If these sets overlap, the procedure continues to the second step; otherwise the players revert to their defaults.
    The intuition for using sets at this stage is that (we will argue) this way a player can use a program that is willing to renegotiate to a payoff above their PMM payoff, without risking miscoordination if the other player does not also choose this new payoff precisely (see Fig. 2). Renegotiation sets that condition on the other player’s renegotiation set function are crucial to the result that players are guaranteed their PMM payoff. This is because unconditionally adding an outcome to the renegotiation set might provide Pareto improvements against some possible counterpart program, but make the outcome worse against some other possible counterpart program (see Example 4.2).
  2. (2)
@@ 4.1. Overview of CSR
4.2. Components of CSR
4.2.1. Set-valued renegotiation
To avoid the need to coordinate on an exact renegotiation function, players can use functions that map miscoordinated outcomes to sets of Pareto improvements they each find acceptable. (In examples, we’ll abuse terminology by referring to action profiles by their corresponding payoff profiles.) Then, we suppose the players follow some rule (a selection function) for choosing an efficient outcome from their agreement set.
@@ 4.2.1. Set-valued renegotiation
Let ${\mathbf{C}}$ $(\mathcal{A})$ be the set of closed subsets of $\mathcal{A}$ [^7] I.e., closed with respect to the topology on $\mathcal{A}$ induced by the Euclidean distance $d(\boldsymbol{a},\boldsymbol{a}^{\prime})=||\textbf{u}(\boldsymbol{a})-\textbf{u}(\boldsymbol{a}^{\prime})||$. Letting $\mathcal{R}^{i}$ be a set of functions from $\mathcal{R}^{j}\times\mathcal{A}$ to $\mathbf{C}(\mathcal{A})$, a function $\mathbf{RN}^{i}\in\mathcal{R}^{i}$ is a set-valued renegotiation function if, for all $\mathbf{RN}^{j}\in\mathcal{R}^{j}$:
  1. (1)
    For all $\boldsymbol{a}\in\mathcal{A}$ and $\boldsymbol{a}^{\prime}\in\mathbf{RN}^{i}(\mathbf{RN}^{j},\boldsymbol{a})$, we have $\textbf{u}(\boldsymbol{a}^{\prime})\succeq\textbf{u}(\boldsymbol{a})$.
  2. (2)
    For some $\boldsymbol{a}$ and some $\boldsymbol{a}^{\prime}\in\mathbf{RN}^{i}(\mathbf{RN}^{j},\boldsymbol{a})$, we have $u_{i^{\prime}}(\boldsymbol{a}^{\prime})>u_{i^{\prime}}(\boldsymbol{a})$ for some $i^{\prime}$.
@@
One might worry that by assuming a fixed selection function, we still haven’t avoided the need for coordination. However, note that there is no bargaining problem involved in coordinating on a selection function. To see this, consider two players who intended to use renegotiation programs with different selection functions. Each player could switch to using a program that used the other player’s selection function, and modify their set-valued renegotiation function so as to guarantee the same outcome as if the other player switched to their selection function. (See Appendix B for a formal argument.) So the players are indifferent as to which selection function is used. (Coordinating on a selection function is a pure coordination problem, however; compare to the problem of coordinating on the programming language used in syntactic comparison-based program equilibrium (Tennenholtz, 2004).) In the results that follow, we will show that players can guarantee the PMM no matter which (transitive) selection function they use.
@@
Figure 2. Set-valued renegotiation in the Scheduling Game, for two possible player 2 renegotiation sets. Black points represent renegotiation outcomes (mapped from the miscoordination outcome $(0,0)$). If player 1 uses the renegotiation set shown here, they can achieve a Pareto improvement even if players don’t reach the Pareto frontier (left), while still allowing for their best possible outcome (right).
[…]
This figure shows two plots illustrating set-valued renegotiation in the Scheduling Game, starting from a miscoordination outcome at (0,0). Both plots show Player 1’s renegotiation set as a shaded region that includes their most preferred outcome (3,1) and all points that are Pareto-worse than (3,1). The left plot shows a case where Player 2’s renegotiation set contains only their most preferred outcome (1,3), resulting in a renegotiation outcome (black point) that is better than (0,0) but not Pareto efficient. The right plot shows a case where Player 2’s set also includes (3,1), allowing the players to achieve Player 1’s most preferred outcome.
@@
(Set-valued renegotiation) Suppose players in the Scheduling Game (Table 1) miscoordinate at $\boldsymbol{a}=(0,0)$. The two plots in Fig. 2 illustrate set-valued renegotiation for two possible player 2 renegotiation sets $\mathbf{RN}^{2}(\mathbf{RN}^{1},\boldsymbol{a})$, and a fixed player 1 renegotiation set $\mathbf{RN}^{1}(\mathbf{RN}^{2},\boldsymbol{a})$. Black points indicate the corresponding renegotiation outcomes. Player 1 thinks it’s likely that the only efficient outcome player 2 is willing to renegotiate to is their own most preferred outcome (1, 3) (topmost gray point, left plot). But player 1 believes that with positive probability player 2’s renegotiation set will also include player 1’s most preferred outcome (3, 1) (black point, right plot). Player 1’s best response given these beliefs may be to choose a set-valued renegotiation function $\mathbf{RN}^{1}$ that maps (0, 0) to a set including both (3, 1) and all outcomes Pareto-worse than (3, 1), i.e., the set depicted in Fig. 2. This way, they still achieve a Pareto improvement if player 2 has the smaller set (left plot), and get their best payoff if player 2 has the larger set (right plot).
4.2.2. Conditional renegotiation sets.
We saw that renegotiation sets allow a player to achieve Pareto improvements against a wider variety of other players than is possible with renegotiation functions. However, suppose a player could not condition their renegotiation set on the other player’s program. Then, by adding a point to their renegotiation set in attempt to Pareto-improve against some possible players, they might lock themselves out of a better outcome against other possible players.
@@
PMM
Figure 3. Two possible renegotiation procedures in the Scheduling Game, for different player 2 renegotiation sets. Player 1 might add the PMM (star) to their unconditional renegotiation set. In the case in the left plot, player 1 is no worse off by adding the PMM to their set. But in the case in the right plot, if player 1 adds the PMM, they might do worse if the selection function chooses the PMM instead of the black point that would have otherwise been achieved.
[…]
This figure shows two plots of renegotiation procedures in the Scheduling Game with different Player 2 renegotiation sets, illustrating why players need conditional renegotiation sets. A star marks the PMM outcome at (1,1). In the left plot, Player 1’s unconditional renegotiation set includes both their original outcome (black point) and the PMM, and they are no worse off when the PMM is chosen. In the right plot, Player 1’s unconditional set again includes both outcomes, but here they would do worse if the selection function chooses the PMM instead of the black point that would have been achieved with their original set.
@@
[…]
4.2.3. Renegotiation sets that guarantee the PMM
How can a player guarantee a payoff better than some miscoordination outcome, without losing the opportunity to bargain for their most-preferred outcome? Suppose player $i$ considers using a set-valued renegotiation function that doesn’t guarantee the PMM. That is, against some counterpart program, the resulting outcome $\boldsymbol{a}$ is worse for at least one player than their least-preferred efficient outcome (i.e., their PMM payoff). Then:
  1. (1)
    As we will argue in Theorem 9, under mild assumptions, player $i$ is no worse off also including $\mathbf{PMP}_{i}(\boldsymbol{a})$ in their renegotiation set. So, if player $j$ also follows the same incentive to include $\mathbf{PMP}_{j}(\boldsymbol{a})$ in their renegotiation set, these programs will guarantee at least the PMM. (Notice that players’ ability to guarantee the PMM depends on conditional renegotiation sets, for the reasons discussed in Example 4.2.)
  2. (2)